Im stuck.
8(3a-5)-4(2a+3)=12
Can you plz help me

Answer by rapaljer(4551)

Watch out for the negative in front of the 4 on this problem. This is the most likely place to make an error:

Remove parentheses by distributive property:
8(3a-5)-4(2a+3)=12

24a -40 -8a - 12 = 12
Combine like terms on the left side:
16a- 52 = 12

Add +52 to each side:
16a - 52+52 = 12 + 52
16a=64

Divide both sides by 16.
a=4

R^2 at SCC